| The Web Newsletter | |
LIGO Livingston Observatory News
Installation Progress at Livingston!
LIGO Summer Student Seminar Series
Behind the Scenes at Livingston
Detector installation continues to make progress in many areas at the LIGO Livingston Observatory (LLO). The seismic isolation work is well underway. Seismic isolation piers, support tubes and tables, air bearing assemblies, and the damped spring assemblies have been installed in Horizontal Access Modules (HAM's) one through four, and in Beam Splitter Chambers (BSC's) one and three. Preparation for installation of the seismic supports in BSC2--the last remaining chamber to be undertaken in the Laser and Vacuum Equipment Area (LVEA)--is underway now. (See Figure 1 below.) Work to complete the installation of major seismic components in the end stations should conclude in October. The air bearing control rack was been installed by Hytech and is now operational. It is connected to BSC3 and will be connected to BSC1 and 2 later this year. Delivery of fine actuators for precision alignment of the interferometer has just begun. They will be installed after all of the BSC work completes.
![]() |
![]() |
Above: Figure 1 at right shows the installation of seismic isolation components in a Horizontal Access Module chamber at the Livingston Observatory. At left in Figure 2 is seen the installation of Beam Tube insulation material.
Preparation for the first Beam Tube bake out at LLO is nearly complete. All of the instrumentation and cabling were shipped from the LIGO Hanford Observatory (LHO) beginning last spring and these components have been re-installed here along the half of the west Beam Tube arm farthest from the corner station. All five miles of Beam Tube were insulated (Figure 2) and that work was completed the beginning of June. The gate valves at the ends and mid-points of the beam were serviced in July by GNB, their manufacturer. The viton O-ring material within the gate valves was replaced with Fluorel, which is more robust at elevated temperatures. Replacement with this material, along with our plan to implement a "soft closure" of the gate valves during bake out (so that the O-ring material is not strongly compressed), should insure that the O-ring material will continue to perform as intended.
The data acquisition and monitoring instrumentation for the bake out is now fully on-line. The Beam Tube is at approximately 10-7 torr and all cryopumps are running. The heater blankets have all been installed and are being tested, as are the main power supplies. First indications are that the summer heat in Louisiana should not prevent the bake out from occurring. We plan to begin the bake in the latter half of August.
Below: Figure 3 at left shows the control racks inside the Laser Vacuum Equipment Area. Next, Figure 4 is an outstanding view of pre-stabilized laser installation in progress.
![]() |
![]() |
The basic data collection software for the interferometer control and data acquisition system is now running. Fabrication and installation of data acquisition and control racks is underway, as shown in Figure 3 at left. Installation of cable trays in the LVEA and end stations was also completed and cable installation is now in progress.
Work on the pre-stabilized laser (Figure 4 at right) is also on-going. The control racks have been installed and the frequency stabilization servo is being implemented.
This summer, a science seminar series was organized
which was aimed specifically at our summer students.
The seminars were ongoing through August and covered a
broad range of scientific topics. Anyone interested was invited to attend,
although students had first priority on asking questions (and
on the pizza which was usually served).
The schedule for August was:
August 3:
"Exploding Stars - Living Fast and Dying Young,"
Professor Bob Svoboda, LSU
August 11:
"An Attitude Determination System for Balloon Astronomy,"
Professor Dick Greenwood, Louisiana Tech University
August 17:
"Cosmology and the Critical Density of the Universe,"
Professor Rai Weiss, MIT
August 18:
"Calculations of Non-linear Optical Effects,"
Professor Bill Parker, Southeastern Louisiana University
August 19:
"Neutrino Physics and Astrophysics,"
Professor Barry Barish, Caltech
August 20:
"Birth of Planetary Systems Directly from Super Novae,"
Dr. Bill Brown, Los Alamos National Laboratory
August 24:
"The Potential for Sapphire Optics in LIGO,"
Professor Steve McGuire, Southern University
August 31:
"Gravity Gradient Measurements at LLO"
Matt Ashman, Caltech
and
"Residual Gas Measurements During the Beam Tube Bake Out"
Quincy Robertson, Southeastern Louisiana University
Matt Ashman (pictured at left) is one of several students working at the
LIGO Livingston Observatory this summer. Matt is a native of this area
(Baton Rouge) and is presently a sophomore at Caltech in Pasadena, California.
Here at Livingston, he has been working on seismic characterization and
gravity gradient noise. And during the school year at Caltech, Matt is
majoring in physics and has been chosen to work and conduct research under
a grant from the Tom Tombrello Physics II Research Tutorial Class. For
personal enjoyment Matt likes playing jazz music. He currently plays string
bass in local clubs and looks forward to finding opportunities around Caltech
and the Los Angeles area to use and expand his abilities.
Quincy Robertson (shown at right) is another one of our summer students. Quincy is a senior and a physics major at Southeastern Louisiana University (SLU). While there Quincy has received several awards for academic excellence and is an active member of the Alpha Phi Alpha fraternity. Here at Livingston, Quincy is working on the Beam Tube bake out project and has been conducting residual gas analysis of the Beam Tube to determine its binding energy. Quincy is not only a student and full time employee but an active father of two boys, Chancellor and Little Quincy.
There is a little stray dog (Rosco, Bubba or Ogil, depending on
who's calling him--not that it really matters because he will
answer to anything) that has made a home here. So far many have
been very generous in caring for the animal. To mention just a
few, Gerry Stapfer has donated a collar and some veterinary care;
Joe Giamie's neighbor (a vet) has donated some samples of a premium
dog food; and Jonathan Kern's boys have donated lots of weekend time
"walking the dog." We interviewed Ogil (Ligo spelled backwards)
for a position as guard dog but we don't think he qualifies. He's far
too friendly--doesn't even bark, much less bite.
We have another animal which has made LIGO it's new home. See it there at right? He doesn't have a name yet. The local hunting club president donated this gator to us. It is believed to be living in the right end station moat, and although several employees have confirmed sightings, I have been unable to catch him on film--so don't feel bad if you didn't see him in the picture, he's not there. Naming contest rules and complete photo to follow in next month's article.